The radius of a can soup is 2 in. and the height is 7 in.
If the radius of the can is cut in half and the height is doubled, what effect will this have on the amount of soup in the can? Provide a mathematical explanation of your answer.

Answer by RAY100(1637)

Remember,, VOLUME is Area of base * Height
.
V1 = pi * r^2 * h
.
V2 = pi *(r/2)^2 * (2h) = pi * r^2 / 4 * 2h = (pi * r^2 *h) * (1/2) = (1/2) * V1
.
numerically,,, V1 = pi * 2^2 * 7 = 97.96 = 88
.
V2 = pi * (2/2)^2 * (2)(7) = pi *1^2 *14 = 43.98 = 44
.